\begin{thebibliography}{HD}

%% Use the widest label as the parameter.
%% Reference items can be numbered or have labels of your choice, as below.

%% In IMPAN journals, only the title is italicized; boldface is not used.
%% Our software will add links to many articles; for this, enclosing volume numbers in { } is helpful
%% Do not give the issue number unless the issues are paginated separately.



%%%%%%% To ease editing, use normal size:

\normalsize
\baselineskip=17pt

%%%%%%%%%%%%%%%
 %Jhorar, Bablesh; Khanduja, Sudesh K. On power basis of a class of algebraic number fields. Int. J. Number Theory 12 (2016), no. 8, 2317–2321.

\bibitem{ANHam} S. Ahmad, T. Nakahara and A. Hameed, {On certain pure sextic fields related to a problem of Hasse}, \emph{Internat. J. Algebra Comput.} {\bf 26} (2016), no. 3, 577--583.

\bibitem{ANHus} S. Ahmad, T. Nakahara and S. M. Husnine, {Power integral bases for certain pure sextic fields}, \emph{Int. J. Number Theory} {\bf 10} (2014), no. 8, 2257--2265.



%\bibitem{AW} \c{S}. Alaca and K. S. Williams, Introductory algebraic number theory, \emph{Cambridge University Press}, Cambridge, 2004.

%\bibitem{Apostol} T. M. Apostol, \emph{Resultants of cyclotomic polynomials}, Proc. Amer. Math. Soc. {\bf 24} (1970) 457--462.
%\bibitem{BB}
%John~A. Beachy and William~D. Blair,
%\emph{Abstract Algebra, third edition}.
%Waveland Press, Inc., 2005.

%\bibitem{BMS1}  B. Bensebaa, A. Movahhedi and A. Salinier, \emph{The Galois group of $x^p+ax^s+a$}, Acta Arith. {\bf 134} (2008), no. 1, 55--65.

%\bibitem{BMS2}   B. Bensebaa, A. Movahhedi and A. Salinier, \emph{The Galois group of $x^p+ax^{p-1}+a$}, J. Number Theory {\bf 129} (2009), no. 4, 824--830.

\bibitem{BSW} M. Bhargava, A. Shankar and X. Wang, {Squarefree values of polynomial discriminants I}, \url{https://arxiv.org/abs/1611.09806}.

%\bibitem{BK}  A. Bishnoi and S. K. Khanduja, \emph{A class of trinomials with Galois group $S_n$}, Algebra Colloq. {\bf 19} (2012), Special Issue no. 1, 905--911.

\bibitem{BMT} D. W. Boyd, G. Martin and M. Thom, {Squarefree values of trinomial discriminants}, \emph{LMS J. Comput. Math.} {\bf 18} (2015), no. 1, 148--169.

%\bibitem{Brandl}  R. Brandl, D. Bubboloni and I. Hupp, \emph{Polynomials with roots mod $p$ for all primes $p$}, J. Group Theory {\bf 4} (2001), no. 2, 233--239.


%\bibitem{BSY}  S. Brown, B. Spearman and Q. Yang, \emph{On sextic trinomials with Galois group $C_6$, $S_3$ or $C_3\times S_3$}, J. Algebra Appl. {\bf 12} (2013), no. 1, 1250128, 9 pp.

%\bibitem{BE}  N. Bruin and N. D. Elkies, \emph{Trinomials $ax^7+bx+c$ and $ax^8+bx+c$ with Galois groups of order 168 and $8\cdot 168$}, Algorithmic number theory (Sydney, 2002), 172--188, Lecture Notes in Comput. Sci., 2369, Springer, Berlin, (2002).

%\bibitem{BM} G. Butler and J. McKay, \emph{The transitive groups of degree up to eleven}, Comm.
%Algebra {\bf 11} (1983), no. 8, 863--911.

\bibitem{Cohen}
H. Cohen, \emph{A Course in Computational Algebraic Number Theory}, Springer-Verlag, 2000.

%\bibitem{CMS} S. D. Cohen, A. Movahhedi and A. Salinier, {Galois groups of trinomials}, \emph{J. Algebra} {\bf 222} (1999), no. 2, 561--573.

\bibitem{Conrad} K. Conrad,\\ \url{www.math.uconn.edu/~kconrad/blurbs/gradnumthy/totram.pdf}

%\bibitem{Conrad2} K. Conrad,\\ \url{www.math.uconn.edu/~kconrad/blurbs/galoistheory/galoisSnAn.pdf}

\bibitem{Cullinan} \url{https://studylib.net/doc/8187082/the-discriminant-of-a-composition-of-two}

%\bibitem{Dietmann} R. Dietmann, \emph{Probabilistic Galois theory}, Bull. Lond. Math. Soc. {\bf 45} (2013), no. 3, 453--462.
\bibitem{Dresden}  G. Dresden, {Resultants of cyclotomic polynomials}, \emph{Rocky Mountain J. Math.} {\bf 42} (2012), no. 5, 1461--1469.
\bibitem{ESW}
D. Eloff, B. Spearman and K. Williams, {$A_4$-sextic fields with a power basis}, \emph{Missouri Journal of Mathematical Sciences} 19 (2007), 188--194.

%\bibitem{Erdos53}
%P. Erd\H{o}s, \emph{Arithmetical properties of polynomials}, The Journal of the London Mathematical Society, \textbf{28} (1953), 416–-425.

%\bibitem{Estermann} T. Estermann, \emph{Einige S\"{a}tze \"{u}ber quadratfreie Zahlen}, Math. Ann., {\bf 105} (1931), 653--662.

\bibitem{GP1}  I. Ga\'{a}l and M. Pohst, \emph{Computing power integral bases in quartic relative extensions}, J. Number Theory {\bf 85} (2000), no. 2, 201--219.

\bibitem{G1} I. Ga\'{a}l, \emph{Power integral bases in cubic relative extensions}, Experiment. Math. {\bf 10} (2001), no. 1, 133--139.

\bibitem{Gaal} I. Ga\'{a}l, \emph{Diophantine equations and power integral bases. New computational methods.}, Birkhäuser Boston, Inc., Boston, MA, (2002).


\bibitem{GR1} I. Ga\'{a}l and L. Remete, {Power integral bases in a family of sextic fields with quadratic subfields}, \emph{Tatra Mt. Math. Publ.} {\bf 64} (2015), 59--66.

\bibitem{GR2} I. Ga\'{a}l and L. Remete, {Integral bases and monogenity of pure fields}, \emph{J. Number Theory} {\bf 173} (2017), 129--146.

\bibitem{GRS}  I. Ga\'{a}l, L. Remete and T. Szab\'{o}, {Calculating power integral bases by using relative power integral bases}, \emph{Funct. Approx. Comment. Math.} {\bf 54} (2016), no. 2, 141--149.

%\bibitem{GS}  I. Ga\'{a}l and T. Szab\'{o}, {Power integral bases in parametric families of biquadratic fields}, \emph{JP J. Algebra Number Theory Appl.} {\bf 24} (2012), no. 1, 105--114.

\bibitem{GS1}  I. Ga\'{a}l and T. Szab\'{o}, \emph{Relative power integral bases in infinite families of quartic extensions of quadratic fields}, JP J. Algebra Number Theory Appl. {\bf 29} (2013), no. 1, 31--43.

\bibitem{Gassert} T. A. Gassert, {Discriminants of Chebyshev radical extensions}, \emph{J. Th\'{e}or. Nombres Bordeaux} {\bf 26} (2014), no. 3, 607--634.

%\bibitem{Granville} A. Granville, \emph{ABC allows us to count squarefrees}, Int. Math. Res. Notices {\bf 1998(19)}
%(1998) 991--1009.

\bibitem{Guerrier} W. J.  Guerrier, {The factorization of the cyclotomic polynomials mod $p$}, \emph{Amer. Math. Monthly} {\bf 75} (1968) 46.


%\bibitem{HJ} J. Harrington and Lenny Jones, \emph{The irreducibility of power compositional sextic polynomials and their Galois groups}, Mathematica Scandinavica (to appear).

%\bibitem{Helfgott} H. A. Helfgott, \emph{On the square-free sieve}, Acta Arith. {\bf 115} (2004), 349--402.

%\bibitem{Helfgott-cubic}  H. A. Helfgott, \emph{Square-free values of $f(p)$, $f$ cubic.}, Acta Math. {\bf 213} (2014), no. 1, 107--135.

%\bibitem{Hooley} C. Hooley, \emph{On the power free values of polynomials}, Mathematika {\bf 14} (1967), 21--26.



%\bibitem{IJ}  J. Ide and Lenny Jones, \emph{Infinite families of $A_4$-sextic polynomials}, Canad. Math. Bull. {\bf 57} (2014), no. 3, 538--545.

\bibitem{Ireland-Rosen} Kenneth Ireland and Michael Rosen, \emph{A Classical Introduction to Modern Number Theory, Second Edition}, {Graduate Texts in Mathematics}, {\bf 84}, Springer--Verlag, New York, 1990.

%\bibitem{JLY} C. U. Jensen, Arne Ledet and Noriko Yui, \emph{Generic polynomials.
%Constructive aspects of the inverse Galois problem}, Mathematical Sciences Research Institute Publications, {\bf 45}, Cambridge University Press, Cambridge, 2002.



%\bibitem{JKS1} A. Jakhar, S. Khanduja and N. Sangwan, \emph{On prime divisors of the index of an algebraic integer}, J. Number Theory {\bf 166} (2016), 47--61.

%\bibitem{JKS2} A. Jakhar, S. Khanduja and N. Sangwan, \emph{Characterization of primes dividing the index of a trinomial}, Int. J. Number Theory {\bf 13} (2017), no. 10, 2505--2514.


\bibitem{JK} B. Jhorar and S. Khanduja, {On power basis of a class of algebraic number fields},
\emph{Int. J. Number Theory} {\bf 12} (2016), no. 8, 2317--2321.

%\bibitem{JP}  L. Jones and T. Phillips, \emph{Infinite families of monogenic trinomials and their Galois groups}, Internat. J. Math. {\bf 29} (2018), no. 5, 1850039, 11 pp.

%\bibitem{K} K. Kedlaya, \emph{A construction of polynomials with squarefree discriminants}, %(English summary)
%Proc. Amer. Math. Soc. {\bf 140} (2012), no. 9, 3025--3033.




%\bibitem{K1}  K. Komatsu, \emph{On the Galois group of $x^p+ax+a=0$}, Tokyo J. Math. {\bf 14} (1991), no. 1, 227--229.

%\bibitem{K2}  K. Komatsu, \emph{On the Galois group of $x^p+p^tb(x+1)=0$}, Tokyo J. Math. {\bf 15} (1992), no. 2, 351--356.

%\bibitem{LS} M. J. Lavallee and B. K. Spearman, \emph{Trinomials $x^6+Ax+B$ defining sextic fields with a pure cubic subfield}, Int. J. Pure Appl. Math. {\bf 56} (2009), no. 3, 345--352.

%\bibitem{MM} G. Malle and B. H. Matzat, \emph{Inverse Galois theory}, Springer-Verlag,
%  Berlin, (1999).

%\bibitem{Mirsky} L. Mirsky, \emph{The number of representations of an integer as the sum of a prime and a $k$-free integer}, Amer. Math. Monthly {\bf 56}, (1949), 17--19.

%\bibitem{M} A. Movahhedi, \emph{Galois group of $x^p+ax+a$}, J. Algebra {\bf 180} (1996), no. 3, 966--975.

%\bibitem{Nagel}
%T. Nagel, \emph{Zur Arithmetik der Polynome}, Abh. Math. Sem. Hamburg. Univ. {\bf 1} (1922), 179--194.

\bibitem{Nagell} T. Nagell, \emph{Introduction to Number Theory}, {New York: Wiley}, 1951.

%\bibitem{N}  W. Narkiewicz, \emph{Elementary and analytic theory of algebraic numbers, Third edition}, Springer Monographs in Mathematics, Springer-Verlag, Berlin, (2004).

%\bibitem{Nivenetal}  I. Niven, H. S. Zuckerman and H. L. Montgomery, \emph{An introduction to the theory of numbers, Fifth edition}, John Wiley \& Sons, Inc., New York, (1991).


%\bibitem{Osada1}  H. Osada, \emph{The Galois groups of the polynomials $x^n+x^l+b$}, J. Number Theory {\bf 25} (1987), no. 2, 230--238.

%\bibitem{Osada2}  H. Osada, \emph{The Galois groups of the polynomials $x^n+ax^s+b$ II}, Tohoku Math. J. (2) {\bf 39} (1987), no. 3, 437--445.

%\bibitem{Pasten} H. Pasten, \emph{The ABC conjecture, arithmetic progressions of primes and squarefree values of polynomials at prime arguments}, Int. J. Number Theory {\bf 11} (2015), no. 3, 721--737.

%\bibitem{Prachar} K. Prachar, \emph{Primzahlverteilung}, Springer, New York, (1957).
%\bibitem{Prachar1}  K. Prachar, \emph{\"{U}ber die kleinste quadratfreie Zahl einer arithmetischen Reihe}, (German) Monatsh. Math. {\bf 62} (1958), 173--176.

%\bibitem{RRS} L. J. Ratliff, D. E. Rush and K. Shah, \emph{Power integral bases for Selmer-like number fields}, J. Number Theory {\bf 121} (2006), no. 1, 90--113.

%\bibitem{Ricci} G. Ricci, \emph{Ricerhe aritmetiche sui polinomi}, Rend. Circ. Mat. Palermo {\bf 57} (1933), 433-475.

%\bibitem{Schinzel}  A. Schinzel, \emph{On reducible trinomials}, Dissertationes Math. (Rozprawy Mat.) {\bf 329} (1993), 83 pp.


%\bibitem{Selmer} E. Selmer, \emph{On the irreducibility of certain trinomials}, Math. Scand. {\bf 4} (1956), 287--302.

%\bibitem{Serre}  Jean-Pierre Serre, \emph{Topics in Galois theory. Second edition. With notes by Henri Darmon.}, Research Notes in Mathematics, {\bf 1.} A K Peters, Ltd., Wellesley, MA, (2008).

%\bibitem{Shpar} I. E. Shparlinski, \emph{Squarefree parts of discriminants of trinomials},
Arch. Math. (Basel) {\bf 102} (2014), no. 6, 545--554.

\bibitem{S1} B. K. Spearman, {Monogenic $A_4$ quartic fields}, \emph{International Mathematical Forum}, {\bf 1}, (2006), no. 40, 1969--1974.

%\bibitem{S2} B. K. Spearman, \emph{Trinomials $x^6+Ax+B$ defining sextic fields with a cyclic cubic subfield}, J. Appl. Algebra Discrete Struct. {\bf 4} (2006), no. 3, 149--155.

%\bibitem{SSW} B. K. Spearman, L. Y. Spearman and K. Williams, \emph{Quadratic subfields of the splitting field of a dihedral quintic trinomial $x^5+ax+b$}, New Zealand J. Math. {\bf 26} (1997), no. 2, 293--299.

%\bibitem{SW1} B. K. Spearman and K. Williams, \emph{Quartic trinomials with Galois groups $A_4$ and $V_4$}, Far East J. Math. Sci. (FJMS) {\bf 2} (2000), no. 5, 665--672.

%\bibitem{SW2} B. K. Spearman and K. Williams, \emph{The discriminant of a dihedral quintic field defined by a trinomial $x^5+ax+b$}, Canad. Math. Bull. {\bf 45} (2002), no. 1, 138--153.

%\bibitem{SW3} B. K. Spearman and K. Williams, \emph{The simplest $D_4$-octics}, Int. J. Algebra {\bf 2} (2008), no. 1--4, 79--89.

%\bibitem{Swan}  R. Swan \emph{Factorization of polynomials over finite fields}, Pacific J. Math. {\bf 12} (1962), 1099--1106.

%\bibitem{U}  K. Uchida, \emph{Unramified extensions of quadratic number fields II}, T\^{o}hoku Math. J. {\bf 2} 22 (1970) 220--224.


%\bibitem{vdWaerden} B. L. van der Waerden, \emph{Die Seltenheit der reduziblen Gleichungen und der Gleichungen mit Affekt}, (German) Monatsh. Math. Phys. {\bf 43} (1936), no. 1, 133--147.

%\bibitem{Walfisz}  A. Walfisz, \emph{Zur additiven Zahlentheorie II}, (German) Math. Z. {\bf 40} (1936), no. 1, 592--607.

%\bibitem{Ward}  M. Ward, \emph{On the factorization of polynomials to a prime modulus}, Ann. of Math. (2) {\bf 36} (1935), no. 4, 870--874.

\bibitem{Washington}  L. C. Washington, \emph{Introduction to cyclotomic fields, Second edition}, {Graduate Texts in Mathematics}, {\bf 83}, Springer-Verlag, New York, 1997. %xiv+487 pp.

% \bibitem{W}  C. J. Williamson, \emph{Odd degree polynomials with dihedral Galois groups},  J. Number Theory {\bf 34} (1990), no. 2, 153--173.

% \bibitem{Y} Y. Yamamoto, \emph{On unramified Galois extensions of quadratic number fields},
%Osaka J. Math. {\bf 7} (1970) 57--76.
\end{thebibliography}

\documentclass[12]{amsart}
\usepackage{amsmath,amssymb,amsthm,color,enumerate,comment,centernot,enumitem,url,cite}
\usepackage{graphicx,relsize}
\usepackage{mathtools}
\usepackage{array}
\newcommand{\ord}{\operatorname{ord}}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}[thm]{Lemma}
\newtheorem{conj}[thm]{Conjecture}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{remark}

\theoremstyle{definition}
    \newtheorem{defn}[thm]{Definition}

\theoremstyle{THM}

\newcommand{\abs}[1]{\left|{#1}\right|}
\def\ds{\displaystyle}

\def\R {{\mathcal R}}
\def\Z {{\mathbb Z}}

\def\Q {{\mathbb Q}}

\def\d {{\rm det}}

\def\F {{\mathbb F}}
\def\R {{\mathcal R}}

\def\Z {{\mathbb Z}}
\def\Q {{\mathbb Q}}

\def\Gal{{\mbox {Gal} }}

\def\red#1 {\textcolor{red}{#1 }}
\def\blue#1 {\textcolor{blue}{#1 }}

\usepackage{array}
\numberwithin{equation}{section}
\def\ds{\displaystyle}
\def\Z {{\mathbb Z}}
\begin{document}

\title[Monogenic Cyclotomic Compositions]{Monogenic Cyclotomic Compositions}


\author{Joshua Harrington}
\address{Department of Mathematics, Cedar Crest College, Allentown, Pennsylvania, USA}
\email[Joshua Harrington]{Joshua.Harrington@cedarcrest.edu}

\author{Lenny Jones}
\address{Professor Emeritus, Department of Mathematics, Shippensburg University, Shippensburg, Pennsylvania 17257, USA}
\email[Lenny~Jones]{lkjone@ship.edu}

\date{\today}

\begin{abstract}
Let $m$ and $n$ be positive integers, and let $p$ be a prime. Let $T(x)=\Phi_{p^m}\left(\Phi_{2^n}(x)\right)$, where $\Phi_k(x)$ is the cyclotomic polynomial of index $k$. In this article, we prove that $T(x)$ is irreducible over $\mathbb Q$ and that
\[\left\{1,\theta,\theta^2,\ldots,\theta^{2^{n-1}p^{m-1}(p-1)-1}\right\}\] is a basis for the ring of integers of $\mathbb Q(\theta)$, where $T(\theta)=0$.
\end{abstract}

\subjclass[2010]{Primary 11R04, Secondary 11R09, 11R32, 12F12}
\keywords{monogenic, irreducible, cyclotomic polynomial, composition}

\maketitle
\section{Introduction}\label{Section:Intro}

In this article, unless stated otherwise, polynomials $f(x)\in \Z[x]$ are assumed to be monic, and when we say $f(x)$ is ``irreducible", we mean irreducible over $\Q$.
Let $K$ be an algebraic number field of degree $n$ over $\Q$. For any $\theta\in K$, we let
$\Delta(\theta):=\Delta\left(1,\theta,\theta^2,\ldots,\theta^{n-1}\right)$ denote the discriminant of $\theta$. Similarly, we let $\Delta(f)$ and $\Delta(K)$ denote the discriminants over $\Q$, respectively, of the polynomial $f(x)$ and the field $K$. If $f(x)$ is irreducible, with $f(\theta)=0$ and $K=\Q(\theta)$, then we have the well-known equation \cite{Cohen}
\begin{equation}\label{Eq:Dis-Dis}
\Delta(f)=\Delta(\theta)=\left[\Z_K:\Z[\theta]\right]^2\Delta(K),
\end{equation}
where $\Z_K$ is the ring of integers of $K$.
We say $f(x)$ is \emph{monogenic} if $f(x)$ is irreducible and $\left[\Z_K:\Z[\theta]\right]=1$, or equivalently from \eqref{Eq:Dis-Dis} that $\Delta(f)=\Delta(K)$. In this situation,
$\left\{1,\theta, \theta^2,\ldots ,\theta^{n-1} \right\}$ is a basis for $\Z_K$ referred to as a \emph{power basis}. The existence of a power basis makes calculations much easier in $\Z_K$. A classic example is the cyclotomic field $K=\Q(\zeta)$, where $\zeta$ is a primitive $n$th root of unity \cite{Washington}. In this case, $\Z_K=\Z[\zeta]$ and so $\Z_K$ has the power basis $\left\{1,\zeta,\ldots,\zeta^{\phi(n)-1}\right\}$, where $\phi(n)$ is Euler's totient function. The minimal polynomial for $\zeta$ over $\Q$ is the cyclotomic polynomial of index $n$, which we denote as $\Phi_n(x)$. That is, $\Phi_n(x)$ is monogenic of degree $\phi(n)$.

Aside from the cyclotomic polynomials, there are many polynomials that are monogenic. In fact, it is shown in \cite{BSW} that the density of the irreducible monogenic polynomials of degree $n$ is $6/\pi^2\approx .607927$. However, finding specific infinite families of monogenic polynomials can be challenging. See for example \cite{ANHam,ANHus,BMT,ESW,Gaal,Gassert,GR1,GR2,JK,S1}.

Motivated by research concerning power bases of relative extensions \cite{GP1,G1,GS1,GRS}, we are led to ask the following basic related question:
\begin{equation}\label{Q:1}
  \mbox{If $f(x)$ and $g(x)$ are monogenic, when is $f\left(g(x)\right)$ monogenic?}
\end{equation}
It is not shocking that a complete answer to \eqref{Q:1} should be difficult to achieve. What is somewhat surprising is that the answer to \eqref{Q:1} requires some effort to unravel, even in seemingly ``easy" situations. For example, although $f(x)=x^2+17$, $g(x)=x^2+5$, $f(g(x))=x^4+10x^2+42$ and $g(f(x))=x^4+34x^2+294$ are all irreducible, and both $f(x)$ and $g(x)$ are monogenic, it turns out that $f(g(x))$ is monogenic but $g(f(x))$ is \emph{not} monogenic.
In this article, we focus our investigation on compositions of certain cyclotomic polynomials. More precisely, we prove the following.
\begin{thm}\label{Thm:Main}
 Let $m$ and $n$ be positive integers, and let $p$ be a prime. Define
 \[T(x):=\Phi_{p^m}\left(\Phi_{2^n}(x)\right),\]
 where $\Phi_k(x)$ is the cyclotomic polynomial of index $k$.
 Then $T(x)$ is irreducible and monogenic.
  \end{thm}

   All computer computations in this article were done using either MAGMA, Maple or Sage.

\section{Preliminaries}\label{Section:Prelim}

We first present some standard terminology and known facts.
\begin{defn}\cite{Cohen}\label{Def:ResDisc}
Let $f(x)$ and $g(x)$ be polynomials over an integral domain $\R$ with respective leading coefficients $a$ and $b$, and respective degrees $m$ and $n$. Let $K$ be the quotient field of $\R$, and let $\overline{K}$ be an algebraic closure of $K$. Suppose that the roots of $f(x)$ and $g(x)$ in $\overline{K}$ are, respectively, $r_1,r_2,\ldots ,r_m$ and $s_1,s_2,\ldots ,s_n$. Then the \emph{resultant} of $f(x)$ and $g(x)$, denoted $R(f,g)$, is defined as
   \[R(f,g):=a^nb^m\prod_{\substack{1\le i\le m\\ 1\le j \le n} }\left(r_i-s_j\right).\] The \emph{discriminant} of $f(x)$, denoted $\Delta(f)$ is defined as
  \[\Delta(f):=\dfrac{(-1)^{m(m-1)/2}}{a}R(f,f^{\prime})=(-1)^{m(m-1)/2}a^{2m-2}\prod_{i\ne j}\left(r_i-r_j\right).\]
 \end{defn}
 \begin{prop}\cite{Cohen}\label{Prop:Res}
   Let $f(x)$ and $g(x)$ be polynomials as in Definition \ref{Def:ResDisc}.
   \begin{enumerate}
   \item $R(f,g)=(-1)^{mn}R(g,f)$,
     \item $R(f,g\cdot h)=R(f,g)R(f,h)$, where $h(x)$ is any polynomial over $\R$,
     \item $\ds R(f,g)=a^n\prod_{1\le i \le m}g\left(r_i\right)$,
     \item \label{I4:Res} Let $c\in \R$. Let $u(x)=f(x+c)$ and $v(x)=g(x+c)$. Then $\ds R(u,v)=R(f,g)$.
          \item \label{I5:Res=0} $f(x)$ and $g(x)$ have a root in common in $\overline{K}$ if and only if $R(f,g)=0$.
   \end{enumerate}
    \end{prop}
    \begin{defn}
  Let $p$ be a prime and let
  \[f(x)=a_nx^n+a_{n-1}x^{n-1}+\cdots +a_1x+a_0\in \Z[x].\] We say $f(x)$ is \emph{$p$-Eisenstein} if
  \[a_n\not \equiv 0 \pmod{p},\quad  a_{0}\equiv a_1\equiv \cdots \equiv a_{n-1}\equiv 0 \pmod{p} \quad \mbox{and} \quad a_0\not \equiv 0 \pmod{p^2}.\]
  \end{defn}

\begin{thm}\cite{Ireland-Rosen} (Eisenstien's Irreducibility Criterion)\label{Thm:Eisenstein's Criterion}
  Let $p$ be a prime and let $f(x)\in \Z[x]$ be $p$-Eisenstein. Then $f(x)$ is irreducible.
\end{thm}

\begin{thm}\label{Thm:Cyclo}
Let $n$ be a positive integer, and let $\Phi_n(x)$ denote the cyclotomic polynomial of index $n$. 
\begin{enumerate}
         \item \label{I:2} \cite{Nagell} Let $p$ be a prime. Then
          \[\Phi_n\left(x^p\right)=\left\{\begin{array}{cl} \Phi_{pn}(x) & \mbox{if $n\equiv 0 \pmod{p}$}\\
      \Phi_n(x)\Phi_{pn}(x) & \mbox{if $n\not \equiv 0 \pmod{p}$}\\
      \end{array}\right.\]
      \item \label{I:3} \cite{Washington} Let $p$ be a prime, and let $m$ be a positive integer. Then
      \[\Delta\left(\Phi_{p^m}(x)\right)=\varepsilon p^{p^{m-1}(pm-m-1)},\]
      \[\mbox{where} \quad \varepsilon=\left\{\begin{array}{rc}
        -1 & \mbox{if $p^m=4$ or $p\equiv 3\pmod{4}$}\\
        1  & \mbox{otherwise}.
      \end{array}\right.\]
      \item \label{I:0} \cite{Guerrier} Let $q$ be a prime such that $q\nmid n$. Let $\ord_n(q)$ denote the order of $q$ modulo $n$. Then $\Phi_n(x)$ factors modulo $q$ into a product of $\phi(n)/\ord_n(q)$ distinct irreducible polynomials, each of degree $\ord_n(q)$. Additionally,
  for any positive integer $m$,
         \[\Phi_{q^mn}(x)\equiv \Phi_n(x)^{\phi\left(q^m\right)} \pmod{q}.\]
         \item \label{Dresden} \cite{Dresden} Let $m$ and $n$ be integers such that $0<m<n$.
             Then
         \[R\left(\Phi_m,\Phi_n\right)=\left\{\begin{array}{cl}
           p^{\phi(m)} & \mbox{if $n/m=p^a$ for some prime $p$ and integer $a\ge 1$,}\\
           1 & \mbox{otherwise.}\\
            \end{array}\right.\]
\end{enumerate}
\end{thm}

The next lemma follows by a straightforward induction on $n$.
\begin{lem}\label{Lem:Induction}
 Let $G(x), H(x)\in \Z[x]$, and let $q$ be a prime. If $G(x)\equiv H(x) \pmod{q}$, then
  \[G(x)^{q^n}\equiv H(x)^{q^n} \pmod{q^{n+1}} \quad \mbox{for all $n\ge 1$}.\]
  \end{lem}

The following theorem gives a formula for the discriminant of the composition of two polynomials. To the best of our knowledge, this result is originally due to John Cullinan \cite{Cullinan} and does not appear in the literature. Hence, for the sake of completeness, we present a proof here.

\begin{thm}\label{Thm:Disccomp}
Let $f(x)$ and $g(x)$ be polynomials in $\mathbb{Q}[x]$, with respective leading coefficients $a$ and $b$, and respective degrees $m$ and $n$.  Then
$$\Delta(f\circ g)=(-1)^{m^2n(n-1)/2}\cdot a^{n-1}b^{m(mn-n-1)}\Delta(f)^{n}R(f\circ g,g')$$
\end{thm}
\begin{proof}
Note that the degree of $f(g(x))$ is $mn$, the leading coefficient of $f(g(x))$ is $ab^m$, the degree of $f^{\prime}(x)$ is $m-1$ and the degree of $f^{\prime}(g(x))$ is $(m-1)n$.
Then
\begin{spreadlines}{.11in}
\begin{align*}
\Delta(f\circ g)
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot R(f\circ g,(f\circ g)')\\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot R(f\circ g,(f'\circ g)\cdot g')\\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot R(f\circ g,f'\circ g)\cdot R(f\circ g,g') \quad \mbox{(Proposition \ref{Prop:Res})} \\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot \left((ab^m)^{(m-1)n}\prod_{\left\{\alpha : f(g(\alpha))=0\right\}}f'(g(\alpha))\right)R(f\circ g,g')\\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot (ab^m)^{(m-1)n}\left(\prod_{\left\{\theta : f(\theta)=0\right\}}f'(\theta)\right)^{n}R(f\circ g,g')\\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot (ab^m)^{(m-1)n}\left(\frac{1}{a^{m-1}}R(f,f')\right)^{n}R(f\circ g,g')\\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot \frac{(ab^m)^{(m-1)n}}{a^{(m-1)n}}R(f,f')^{n}R(f\circ g,g')\quad \mbox{(Proposition \ref{Prop:Res})}\\
=&\frac{(-1)^{mn(mn-1)/2}}{ab^m}\cdot \frac{(ab^m)^{(m-1)n}}{a^{(m-1)n}}\left(\frac{a\Delta(f)}{(-1)^{m(m-1)/2}}\right)^{n}R(f\circ g,g')\\
=&(-1)^{m^2n(n-1)/2}\cdot a^{n-1}b^{m(mn-n-1)}\Delta(f)^{n}R(f\circ g,g').\qedhere
\end{align*}
\end{spreadlines}
\end{proof}

The following special case of Theorem \ref{Thm:Disccomp} will be useful to us.
\begin{cor}\label{Cor:Disccomp}
Let $f(x)\in \mathbb{Q}[x]$, such that $a$ is the leading coefficient of $f(x)$,  and $m=\deg(f)$. If $g(x)=bx^n+c\in \Q[x]$, then
$$\Delta(f\circ g)=(-1)^{mn(n-1)(m+2n)/2} a^{n-1}b^{m(mn-1)}\Delta(f)^{n}n^{mn}f(c)^{n-1}.$$
\end{cor}

\begin{proof} Note that $\deg(f\circ g)=mn$ and $\deg(g')=n-1$. Then, using Theorem \ref{Thm:Disccomp} and Proposition \ref{Prop:Res} we have
\begin{spreadlines}{.11in}
 \begin{align*}
\Delta(f\circ g)
=&(-1)^{m^2n(n-1)/2} a^{n-1}b^{(m-1)nm-m}\Delta(f)^{n}R(f\circ g,g')\\
=&(-1)^{m^2n(n-1)/2} a^{n-1}b^{(m-1)nm-m}\Delta(f)^{n}(-1)^{mn(n-1)}R(g',f\circ g)\\
=&(-1)^{m^2n(n-1)/2} a^{n-1}b^{(m-1)nm-m}\Delta(f)^{n}(-1)^{mn(n-1)}(bn)^{mn}\prod_{i=1}^{n-1}f(g(0))\\
=&(-1)^{m^2n(n-1)/2} a^{n-1}b^{(m-1)nm-m}\Delta(f)^{n}(-1)^{mn(n-1)}(bn)^{mn}f(c)^{n-1}\\
=&(-1)^{mn(n-1)(m+2n)/2} a^{n-1}b^{m(mn-1)}\Delta(f)^{n}n^{mn}f(c)^{n-1}.\qedhere
\end{align*}
\end{spreadlines}
\end{proof}

\begin{thm}\cite{Conrad}\label{Thm:Conrad}
 Let $p$ be a prime and let $f(x)\in \Z[x]$ be a monic $p$-Eisenstien polynomial with $\deg(f)=n$. Let $K=\Q(\theta)$, where $f(\theta)=0$. If $n\not \equiv 0 \pmod{p}$, then $p^{n-1} || \Delta(K)$.
\end{thm}

\begin{thm}[Dedekind] \label{Thm:Dedekind}\cite{Cohen}
Let $K=\Q(\theta)$ be a number field, $T(x)\in \Z[x]$ the monic minimal polynomial of $\theta$, and $\Z_K$ the ring of integers of $K$. Let $p$ be a prime number and let $\overline{ * }$ denote reduction of $*$ modulo $p$ (in $\Z$, $\Z[x]$ or $\Z[\theta]$). Let
\[\overline{T}(x)=\prod_{i=1}^k\overline{t_i}(x)^{e_i}\]
be the factorization of $T(x)$ modulo $p$ in $\F_p[x]$, and set
\[g(x)=\prod_{i=1}^kt_i(x),\]
where the $t_i(x)\in \Z[x]$ are arbitrary monic lifts of the $\overline{t_i}(x)$. Let $h(x)\in \Z[x]$ be a monic lift of $\overline{T}(x)/\overline{g}(x)$ and set
\[F(x)=\dfrac{g(x)h(x)-T(x)}{p}\in \Z[x].\]
Then
\[\left[\Z_K:\Z[\theta]\right]\not \equiv 0 \pmod{p} \Longleftrightarrow \gcd\left(\overline{F},\overline{g},\overline{h}\right)=1 \mbox{ in } \F_p[x].\]
\end{thm}

\subsection{The Proof of Theorem \ref{Thm:Main}}
\begin{proof}
Note that when $n=1$, the theorem is trivially true since in that case $T(x)$ is simply a translation of $\Phi_{p^m}(x)$. So, assume that $n\ge 2$.
 We first show that $T(x)$ is irreducible by showing that $T(x)$ is $p$-Eisenstein. Since $\Phi_1(x)=x-1$ and $\Phi_{2^n}(x)=x^{2^{n-1}}+1$, we have from part \eqref{I:0} of Theorem~\ref{Thm:Cyclo} with $q=p$ and $n=1$ that
\begin{equation}\label{Eq:T mod p}
T(x)=\Phi_{p^m}\left(\Phi_{2^n}(x)\right) \equiv \Phi_1(\Phi_{2^n}(x))^{p^{m-1}(p-1)}\equiv x^{2^{n-1}p^{m-1}(p-1)} \pmod{p}.
\end{equation}
It follows that only the leading coefficient of $T(x)$ is not divisible by $p$.  Since the constant term of $T(x)$ is $T(0)=\Phi_{p^m}(\Phi_{2^n}(0))=\Phi_{p^m}(1)=p$, we conclude that $T(x)$ is $p$-Eisenstein.

Next, using Corollary \ref{Cor:Disccomp} and part \eqref{I:3} of Theorem \ref{Thm:Cyclo}, we have that
\begin{align}\label{Eq:Disc(T)}
\begin{split}
\Delta(T)&=\left(\varepsilon p^{p^{m-1}(pm-m-1)}\right)^{2^{n-1}}\left(2^{n-1}\right)^{p^{m-1}(p-1)2^{n-1}}\Phi_{p^m}(1)^{2^{n-1}-1}\\
&=\varepsilon^{2^{n-1}}2^{(n-1)2^{n-1}p^{m-1}(p-1)}p^{m2^{n-1}p^{m-1}(p-1)-1}.\\
\end{split}
\end{align}
Suppose that $T(\theta)=0$, and let $K=\Q(\theta)$.
We now use Theorem \ref{Thm:Dedekind} to show that neither 2 nor $p$ divides $\left[\Z_K:\Z[\theta]\right]$.

We address the prime $p$ first. From \eqref{Eq:T mod p}, we can let $g(x)=x$ and $g(x)h(x)=x^{2^{n-1}p^{m-1}(p-1)}$ in Theorem \ref{Thm:Dedekind}, so that
\[F(x)=\dfrac{x^{2^{n-1}p^{m-1}(p-1)}-T(x)}{p}.\] It follows that $\overline{F}(0)=-1$. Consequently, $\gcd\left(\overline{F},\overline{g}\right)=1$ in $\F_p[x]$, and $\left[\Z_K:\Z[\theta]\right]\not \equiv 0 \pmod{p}$ by Theorem \ref{Thm:Dedekind}.

If $p=2$, there is nothing else to do. So, assume that $p\ne 2$. To address the prime 2, first note that
\begin{equation*}\label{Eq:Mod2}
   T(x)\equiv \Phi_{p^m}\left((x+1)^{2^{n-1}}\right)\equiv \left(\Phi_{p^m}\left(x+1\right)\right)^{2^{n-1}} \pmod{2}.
\end{equation*}
Let $a=\phi\left(p^m\right)$ and $b=\ord_{p^m}(2)$. By part \eqref{I:0} of Theorem \ref{Thm:Cyclo} with $n=p^m$ and $q=2$, we conclude that
\begin{equation}\label{Eq:Phicong}
\Phi_{p^m}\left(x+1\right)\equiv\prod_{i=1}^{a/b}t_i(x)\pmod{2},
\end{equation}
where the $t_i(x)$ are distinct irreducible polynomials each of degree $b$. Hence,
we may take
\begin{equation}\label{Eq:g}
g(x)=\prod_{i=1}^{a/b}t_i(x)\quad \mbox{and} \quad g(x)h(x)=\left(\prod_{i=1}^{a/b}t_i(x)\right)^{2^{n-1}}
\end{equation} in Theorem \ref{Thm:Dedekind}.
Then, from \eqref{Eq:Phicong}, we can write
\begin{equation}\label{Eq:g exp}
g(x)=\Phi_{p^m}\left(x+1\right)+2r(x)=\frac{(x+1)^{p^m}-1}{(x+1)^{p^{m-1}}-1}+2r(x),
\end{equation}
for some polynomial $r(x)\in \Z[x]$. For $n\ge 2$, define
\begin{equation}\label{Eq:Fn}
F_n(x):=\dfrac{g(x)^{2^{n-1}}-\Phi_{p^m}\left(x^{2^{n-1}}+1\right)}{2} \in \Z[x].
\end{equation} By Lemma \ref{Lem:Induction}, we deduce that $g(x)^{2^n}\equiv g\left(x^2\right)^{2^{n-1}} \pmod{4}$ for $n\ge 2$ so that
\[g(x)^{2^n}-\Phi_{p^m}\left(x^{2^n}+1\right)\equiv g\left(x^2\right)^{2^{n-1}}-\Phi_{p^m}\left(\left(x^2\right)^{2^{n-1}}+1\right)  \pmod{4}.\] That is,
\[2F_{n+1}(x)\equiv 2F_n\left(x^2\right) \pmod{4} \quad \mbox{for $n\ge 2$}.\]
Hence, for all $n\ge 2$, it follows that
\[F_{n+1}(x)\equiv F_n\left(x^2\right)\equiv F_n(x)^2 \pmod{2}.\] Consequently, to show that $\gcd\left(\overline{F_n},\overline{g}\right)=1$ for all $n\ge 2$, it is enough to show that $\gcd\left(\overline{F_2},\overline{g}\right)=1$.
Let
\[A(x)=(x+1)^{p^m}-1, \quad B(x)=\left(x^{2}+1\right)^{p^{m-1}}-1,\]
 \[\quad C(x)=\left(x^2+1\right)^{p^m}-1 \quad \mbox{and} \quad D(x)=(x+1)^{p^{m-1}}-1.\] Using \eqref{Eq:g exp} and \eqref{Eq:Fn}, we have that
\begin{align}\label{Eq:F21}
F_2(x)&=\dfrac{\left(\dfrac{A(x)}{D(x)}+2r(x)\right)^{2}-\dfrac{C(x)}{B(x)}}{2} \nonumber\\
&=\dfrac{A(x)^2B(x)-C(x)D(x)^2}{2B(x)D(x)^2}+2\left(\dfrac{r(x)A(x)+r(x)^2D(x)}{D(x)}\right).
\end{align}
Expanding $A(x)^2$, we get
\begin{align*}
  A(x)^2&=(x+1)^{2p^m}-2(x+1)^{p^m}+1\\
  &=\left(x^2+1+2x\right)^{p^m}-2(x+1)^{p^m}+1\\
  &=\left(\left(x^2+1\right)^{p^m}+p^m\left(x^2+1\right)^{p^m-1}(2x)+\cdots +(2x)^{p^m}\right)-2(x+1)^{p^m}+1\\
  &=\left(x^2+1\right)^{p^m}+2p^mx\left(x^2+1\right)^{p^m-1}+4u(x)-2(x+1)^{p^m}+1,
\end{align*}
where
\[4u(x)=\binom{p^m}{2}\left(x^2+1\right)^{p^m-2}(2x)^2+\cdots +\binom{p^m}{p^m-1}\left(x^2+1\right)(2x)^{p^m-1}+(2x)^{p^m}.\]
Similarly,
\[D(x)^2=\left(x^2+1\right)^{p^{m-1}}+2p^{m-1}x\left(x^2+1\right)^{p^{m-1}-1}+4v(x)-2(x+1)^{p^{m-1}}+1,\]
where 
\[4v(x)=\binom{p^{m-1}}{2}\left(x^2+1\right)^{p^{m-1}-2}(2x)^2+\cdots +(2x)^{p^{m-1}}.\]
Substituting the expressions above for $A(x)^2$ and $D(x)^2$ into the numerator of the first term of \eqref{Eq:F21} and cancelling the factor of 2 yields
\begin{equation*}\label{Eq:F22}
  F_2(x)=\dfrac{E(x)}{B(x)D(x)^2}+2\left(\dfrac{u(x)B(x)-v(x)C(x)}{B(x)D(x)^2}+\dfrac{r(x)A(x)+r(x)^2D(x)}{D(x)}\right),
\end{equation*}
where
\begin{multline*}\label{Eq:E}
E(x)=-\left(x^2+1\right)^{p^m}+\left(x^2+1\right)^{p^{m-1}}+p^mx\left(x^2+1\right)^{p^m-1+p^{m-1}}\\
-p^mx\left(x^2+1\right)^{p^m-1}-(x+1)^{p^m}\left(x^2+1\right)^{p^{m-1}}\\
+(x+1)^{p^m}-\left(x^2+1\right)^{p^m-1+p^{m-1}}p^{m-1}x\\
+\left(x^2+1\right)^{p^m}(x+1)^{p^{m-1}}\\
+p^{m-1}x\left(x^2+1\right)^{p^{m-1}-1}-(x+1)^{p^{m-1}}.
\end{multline*}
Then, reduction of $F_2(x)$ modulo 2 produces
\begin{equation}\label{Eq:F22 mod 2}
  \overline{F_2}(x)=\dfrac{\overline{E}(x)}{\left((x+1)^{p^{m-1}}-1\right)^4},
\end{equation}
where
\begin{multline*}\label{Eq:Ebar}
\overline{E}(x)=\left(x+1\right)^{2p^m}+\left(x+1\right)^{2p^{m-1}}+x\left(x+1\right)^{2\left(p^m-1\right)}+(x+1)^{p^m}\left(x+1\right)^{2p^{m-1}}\\
+(x+1)^{p^m}+\left(x+1\right)^{2p^m}(x+1)^{p^{m-1}}
+x\left(x+1\right)^{2\left(p^{m-1}-1\right)}+(x+1)^{p^{m-1}}.
\end{multline*}
  We claim that $\gcd\left(\overline{F_2},\overline{g}\right)=1$. To establish this claim, assume to the contrary that there exists $\alpha$ in some algebraic closure of $\F_2$ such that
  \begin{equation}\label{Eq:Assumption}
  \overline{g}(\alpha)=0=\overline{F_2}(\alpha).
  \end{equation} Then, from \eqref{Eq:Phicong} and \eqref{Eq:g}, we have $\Phi_{p^m}(\alpha+1)=0$ so that $\left(\alpha+1\right)^{p^m}=1$.
    Hence, it follows from \eqref{Eq:F22 mod 2} that
\begin{spreadlines}{.11in}
\begin{align*}
 \overline{F_2}(\alpha)&= \dfrac{\alpha(\alpha+1)^{2\left(p^{m-1}-1\right)}\left(\left((\alpha+1)^{p^{m-1}}\right)^{(p-1)/2}-1\right)^4}{\left((\alpha+1)^{p^{m-1}}-1\right)^4}\\
&=\alpha(\alpha+1)^{2\left(p^{m-1}-1\right)}\left(\prod_{\substack{d\mid \frac{p-1}{2}\\d>1}}\Phi_d\left((\alpha+1)^{p^{m-1}}\right)\right)^4.
\end{align*}
\end{spreadlines}
Thus, since $\overline{F_2}(\alpha)=0$ and both $\overline{g}(0)$ and $\overline{g}(-1)$ are nonzero, we conclude that $\Phi_{p^m}(x+1)$ and $\Phi_{d}\left((x+1)^{p^{m-1}}\right)$ have the root $\alpha$ in common for some divisor $d>1$ of $(p-1)/2$. Repeated application of part \eqref{I:2} of Theorem \ref{Thm:Cyclo} yields
\begin{equation*}
 \Phi_{d}\left((x+1)^{p^{m-1}}\right)=\prod_{j=0}^{m-1}\Phi_{p^jd}(x+1).
\end{equation*}
Therefore, $\Phi_{p^m}(x+1)$ and $\Phi_{p^jd}(x+1)$ have the root $\alpha$ in common for some $j$ with $0\le j\le m-1$. However, recall that $d$ is a divisor of $(p-1)/2$ with $d>1$ and note that $p^jd<p^m$. Hence, $\gcd(d,p)=1$ and $p^m/p^jd$ is not an integer. Thus, from part \eqref{Dresden} of Theorem \ref{Thm:Cyclo}, we have that $R\left(\Phi_{p^jd},\Phi_{p^m}\right)=1$, which contradicts \eqref{Eq:Assumption} by part \eqref{I5:Res=0} of Proposition \ref{Prop:Res}. Therefore, by Theorem \ref{Thm:Dedekind}, we have established that $\left[\Z_K:\Z[\theta]\right]\not \equiv 0 \pmod{2}$, which completes the proof of the theorem.
\end{proof}

\section{Final Remarks}
We certainly do not claim that Theorem \ref{Thm:Main} represents the most general result possible, since computationally there appear to be many other situations where the composition of two cyclotomic polynomials is indeed monogenic. The main difficulty in establishing a result that is more general than Theorem \ref{Thm:Main} arises from the handling of the prime divisors of the resultant appearing in the formula given in Theorem \ref{Thm:Disccomp} for the discriminant of the composition. On the other hand, while many monogenic cyclotomic compositions exist that are not described by Theorem \ref{Thm:Main}, we should point out that slight deviations from the polynomials in Theorem \ref{Thm:Main} yield compositions that are not monogenic. For example, it is not difficult to show that $\Phi_4\left(\Phi_{3^n}(x)\right)$ and $\Phi_3\left(\Phi_{5^n}(x)\right)$ are reducible for all $n\ge 1$, and although $\Phi_2\left(\Phi_{25}(x)\right)$ is irreducible, it is not monogenic since the ring of integers does not possess a power basis. Perhaps these exceptions can be completely determined, but that is a topic for future research.

\end{document}